OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = -x6 + x5- x4 + 4x3 - 12x2 + 12

OpenStudy (anonymous):

a Positive: 4, 2, or 0; Negative: 2 or 0; Complex: 6, 4, 2, or 0 b Positive: 3 or 1; Negative: 3 or 1; Complex: 4, 2, or 0 c Positive: 5, 3, or 1; Negative: 1; Complex: 4, 2, or 0 d Positive: 2 or 0; Negative: 2 or 0; Complex: 6, 4, or 2

OpenStudy (anonymous):

oik here we go with descartes rule of sign do you know how to use it?

OpenStudy (anonymous):

i think.. would it be option c?

OpenStudy (anonymous):

\[f(x)=x^5- x^4 + 4x^3 - 12x^2 + 12\] has 5 changes in sign, so either 5, 3, or 1 positive zero yes that part is right

OpenStudy (anonymous):

now we need \(f(-x)=-x^5-x^4-4x^3-12x^2+12\) which has only one change in sign

OpenStudy (anonymous):

since you count down by 2's there must be one negative zero

OpenStudy (anonymous):

option.. d?

OpenStudy (anonymous):

ooh damn damn damn!

OpenStudy (anonymous):

i didn't see the first term, have to start again

OpenStudy (anonymous):

\(f(x)= f(x) = -x^6 + x^5- x^4 + 4x^3 - 12x^2 + 12\)

OpenStudy (anonymous):

ok yes 5 changes in sign right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 5,3,1 for positive zeros

OpenStudy (anonymous):

ooohh, so C right?

OpenStudy (anonymous):

\[ f(-x) = -x^6 - x^5- x^4 - 4x^3 - 12x^2 + 12\]

OpenStudy (anonymous):

yes it is C

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw, but i think you did all the work